# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v2

@@ Abstract
Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles. This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph Level Order) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions. By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist. This approach not only confirms the conjecture for all oriented graphs but also provides a novel framework for analyzing degrees and arcs in complex networks. The findings have implications for theoretical graph studies and practical applications in network optimization and algorithm design.
@@ 1 Introduction
Seymour’s Second Neighborhood Conjecture, proposed by Paul Seymour in 1990, seeks to identify a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [11].
@@
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]) to encompass a broader class of oriented graphs, thereby providing a framework applicable across various minimum degree scenarios.
The Second Neighborhood Conjecture presents a compelling challenge within the context of graph theory, focusing on oriented graphs and the behavior of second neighbors in the graph’s square. Efforts to resolve this conjecture reflect the deep intrigue and mathematical beauty inherent in such problems. Beyond its theoretical importance, the conjecture has potential applications in network theory, algorithm design, and combinatorial optimization—fields where optimizing connections between nodes is crucial.
While Seymour’s Conjecture remains an open problem, various efforts have tackled special cases. For example, Kaneko and Locke [12] demonstrated its truth for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square, concluding that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
The problem of identifying structural properties in oriented graphs has been studied extensively, with particular focus on $m$-free graphs and Seymour vertices. Daamouch’s [7] work has been instrumental in this area, introducing significant results regarding induced subgraphs and their role in determining key vertex properties. This paper builds on these foundational ideas while presenting a novel approach that extends these concepts to interconnected container structures.
In this paper, we will prove Seymour’s Conjecture by assuming it is false and then attempting to construct a counterexample. Specifically, we explore the conditions under which no node’s out-degree doubles in the square of an oriented graph. Our approach focuses initially on minimum out-degree nodes, analyzing how their out-degree could fail to double, and gradually extending this analysis to other nodes based on their distance from the minimum out-degree node.
The Seymour Second Neighborhood Conjecture has long been considered an intractable problem in graph theory, with some experts even doubting its validity. This paper presents a rigorous proof of the conjecture, accompanied by extensive supporting materials, including appendices, detailing definitions, lemmas, examples and an algorithm, as well as an interactive website for visualizing key ideas. These resources aim to make the proof as transparent and accessible as possible, acknowledging the high standards of scrutiny such results rightly deserve.
@@ 2 Graph Theory Terminology
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$, which are nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$–that is, nodes that are two steps away from $v$ in $G$.
@@ 3 Approach and Motivation
We begin our analysis by considering general oriented graphs without making any initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs, such as the existence of an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we approach the problem from two perspectives. The first is the graph $k$-partition problem [10], where the original oriented graph $G$ is partitioned into $d_{v_{0}}+1$ subgraphs, corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm, which is widely recognized in computer science for simplifying complex problems by tackling smaller, independent subproblems. These partitions enable us to decouple the analysis of interior relationships within each partition from the exterior interactions between partitions, allowing us to address each separately before examining how they integrate within the overall structure. Although the partitions remain connected through certain edges, each can be analyzed independently, allowing us to address subproblems in isolation before combining results to solve the overall problem.
By analyzing each partition individually, we aim to gain localized insights into the graph’s structure and degree dynamics, which we can then synthesize to form a cohesive solution. This divide-and-conquer approach has demonstrated considerable success in algorithms such as mergesort, quicksort, and matrix multiplication. While divide-and-conquer is not commonly applied in graph theory—largely because graphs are typically viewed as unordered collections of vertices and edges—we seek to view this differently by introducing a total ordering on the nodes of $G$. This will allow us to treat the graph in a way that is amenable to divide-and-conquer strategies.
While previous approaches to Seymour’s Second Neighborhood Conjecture often assume the location of a degree-doubling node and focus their analysis around this fixed point, our approach adopts a search-based strategy. Rather than assuming where the doubling will occur, this paper introduces an algorithm that systematically searches the graph, identifying degree-doubling nodes based on structural properties. This method offers a flexible and robust framework that does not depend on preexisting assumptions, allowing for a comprehensive analysis applicable to a broader range of oriented graphs.
This fundamentally shifts the way we approach graph theory problems by introducing a hierarchical structure that is conducive to divide-and-conquer techniques. This reimagining of graph structure allows us to efficiently isolate subproblems, analyze them independently, and then integrate results—offering both computational efficiency and scalability.
One of the key advantages of the structure is its capacity to handle multiple metrics: an outer metric (e.g., distance) for partitioning the graph, and an inner metric (e.g., degree) for analyzing each partition’s internal structure. This dual-metric approach is powerful because it allows for a more nuanced exploration of graph properties, particularly in cases where traditional methods would be insufficient.
The algorithm introduced in this paper systematically searches the graph, leveraging the graph’s structure, total ordering and degree metrics. By not presupposing the location of degree-doubling nodes, the algorithm dynamically identifies these nodes, ensuring a comprehensive exploration of the graph’s structure.
Previous work on Seymour’s Second Neighborhood Conjecture often centers around assuming the position of a degree-doubling node and anchoring the analysis around this fixed point. In contrast, our approach eliminates this assumption, offering a more flexible framework that dynamically adapts to the structure of the graph. This ability to apply a divide-and-conquer strategy to a broader set of graphs marks a significant step forward in graph theory.
@@ 4 Initial Lemmas
We begin our investigation of oriented graphs by searching for a node whose degree could potentially double. While many candidates exist, our goal is to find a node such that, if its degree does not double, we still gain useful information about the structure of the graph. This approach allows us to proceed iteratively, analyzing the graph node by node until we either exhaust all possibilities or discover a counterexample.
@@
Minimum Out-Degree ¡ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
@@
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0, and hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. Since $v_{0}$ is a minimum out-degree node, we know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$. Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$, meaning $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
@@
Lemma 4.1 Minimum Out-Degree ¡ 3 gives us a glimpse on the picture of an oriented graphs. It tells us that, even with just the concept of a minimum degree node we are able to impose restrictions on not only that node $v_{0}$, but also on every other node $v\in G$. This later part is true because $v_{0}$ is a minimum degree node and so every other node must have at least the out-degree of $v_{0}$. What this lemma is saying is that cases where $v_{0}$ has very few out-neighbors, its second neighborhood will expand significantly. This aligns with the idea that low out-degree nodes should see a substantial increase in neighbors when considering the second neighborhood.
@@
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$ Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
@@
Thus, $v_{0}$’s out-degree will at least double in $G^{2}$.
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors is somewhat ”isolated” within the neighborhood, leading to an increased second neighborhood for $v_{0}$.
@@
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree ¡ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
@@
This lemma underscores a crucial insight: while a node’s out-degree doubles, it is not necessarily the initial minimum out-degree node $v_{0}$ whose degree doubles. Instead, the doubling occurs in one of its neighbors. This distinction is significant because it challenges prior approaches that focus on selecting the ”right” starting node for analysis.
Our approach does not depend on starting with a specific node but instead leverages a search process that inherently reveals the doubling behavior in the graph structure. This framework allows for a more general and robust analysis.
For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph. The graph can be illustrated as follows:
The initial lemmas serve as essential building blocks in our exploration. By establishing these results, we create a solid foundation for the more profound theoretical analysis that follows. Each lemma reinforces the central theme of understanding the interplay between degrees and connectivity in oriented graphs.
@@ 5 Theoretical Foundation
Building on the initial lemmas, this section develops the theoretical foundation necessary for addressing the conjecture. We explore key properties such as containers, the Decreasing Sequence Property and the concepts of interior and exterior degrees, which are vital to our argument.
@@
This definition of a container is conceptually related to Daamouch’s use of induced subgraphs in his study of $m$-free graphs [7]. However, the key difference lies in the formation of connected containers (see image), or neighboring containers, which allows for a more dynamic and scalable analysis of graph properties. This approach facilitates the identification of structural patterns across broader graph classes, a task that would otherwise be computationally prohibitive.
As the previous definition aligns with the concept of double arrows, the term ”containers” allows us to explore the structure of oriented graphs and their interactions with minimum degree nodes and distances from these nodes. By introducing containers, we can focus not only on the properties of the nodes relative to a minimum degree node, but also on their interactions within and across containers.
This framework enables us to examine node behavior in three key aspects. First, we can study node behavior relative to a minimum degree node. Containers group nodes based on their distance from a minimum degree node $v_{0}$, making it easier to analyze how these nodes behave and how their properties evolve as we move further from $v_{0}$.
Second, we can investigate internal interactions within containers. Once nodes are grouped into containers, we can study how nodes within the same container influence one another. These internal interactions may reveal patterns related to degree growth, connectivity, and second-degree effects.
Containers are an attempt to bring order to an oriented graph. We mean that in the literal sense of the word, order. For nodes belonging to two different containers, they can now be compared by their distance to a minimum degree node $v_{0}$ which we call in container 0. Since distance is a total ordering, this allows to compare over every other node in $G$
The next few chapters will speak about how these containers communicate with each other and within themselves. No nodes or arcs are removed from the graph to create these containers. This is just a way of thinking about the graph so as to perform our analysis in search of a counterexample.
Finally, we can explore inter-container interactions. By analyzing interactions between nodes from different containers, we gain insight into how graph components expand or contract, how nodes influence one another across distances, and how arc directions impact the convergence of the algorithm or counterexample we aim to develop.
@@
In previous sections, we discussed two fundamental concepts: the minimum degree node and the decreasing sequence property (DSP). A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point for connectivity and flow within the graph. As the graph’s structural anchor, the minimum degree node drives the development of relationships and properties that will emerge in the counterexample construction. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity by systematically limiting potential connections beyond nodes that exhibit this property. This concept is akin to a bottleneck: even nodes with multiple connections may still encounter restricted reachability in the graph. A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP.
Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively. They jointly reveal scenarios where expected connectivity may falter, highlighting vulnerabilities that could challenge established conjectures. The DSP, in particular, illustrates how exterior nodes in an oriented graph may systematically decrease in number, creating a structure where connectivity flow is constrained despite high out-degrees of certain nodes.
By examining the DSP, we underscore a critical aspect of connectivity, resembling bottlenecks in network theory. Diestel’s Graph Theory [6] addresses the relationship between node degree and connectivity, demonstrating that a node $x$ can maintain many outgoing neighbors while still limiting reachability. This discrepancy signals inefficiencies and sets the stage for potential counterexamples to existing conjectures, thus motivating our investigation of graph properties and their broader implications.
@@
Unlike a simple total order, which could be any sequence with a single, consistent ordering (like lexicographic ordering), the container structure allows nodes at the same distance to be further organized. This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties. This layered organization is especially useful in proofs, where we might need to argue about the relative positions of nodes within the same distance from $v_{0}$.
This structure allows us to establish a refined form of comparability among nodes at equal distance that is crucial for analyzing specific properties. For example, in certain proofs, we might need to show that nodes within the same container have certain relationships that would not necessarily follow from a mere total order.

### v3

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods. This data structure allows us to construct a decreasing sequence of subsets, ensuring that no counterexample can exist. Further, this algorithmic approach allows us to know where the node is. Later findings extend the results to multiple nodes. This is true for all oriented graphs. These findings also have practical applications in network optimization and algorithm design.
@@ 1 Introduction
Seymour’s Second Neighborhood Conjecture (SSNC) was originally proposed by Paul Seymour in 1990. It asks the question of the existence of a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [10].
@@
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]). Here, we hope to tackle the entire class of oriented graphs based on the structure laid out in that paper.
The Seymour Second Neighborhood Challenge is a difficult graph theory problem. The problem itself asks us to focus on the question of oriented graphs and second neighbors. There have been several papers that have looked into special cases. Kaneko and Locke [11] showed that the conjecture was true for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square. They concluded that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here. In his paper, he introduces induced subgraphs and speaks of their roles in degree properties on vertices, which will also be a key thing in this paper.
The Seymour Second Neighborhood Conjecture is not strictly a graph theory problem. Although it has historically been studied within the field of mathematics, it has practical applications across several domains, including social media, social networking, epidemiology, disease control, and many other fields. In the critical context of disease control, the nodes in this conjecture can represent individuals, and the orientation of the arcs symbolizes one person infecting another. Here, the application of SSNC becomes relevant by asking: Is there a person whose influence in their second radius of interaction (i.e., people they indirectly connect with) exceeds the number of people in their first radius (i.e., direct contacts)?
In this paper, we will prove the Seymour Second Neighborhood Conjecture is true. We will do this by first assuming that it is false and attempting to construct a counterexample. We will explore the conditions under which no node’s out-degree doubles will double in the square of the oriented graph. This approach will focus initially on minimum out-degree nodes. We will analyze how their out-degree could fail to double. We will then extend out algorithm to other nodes based on their distance from the minimum out-degree node.
@@ 2 Graph Theory Terminology
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$. These are the nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$ that is, nodes that are two steps away from $v$ in $G$.
@@ 3 Approach and Motivation
We begin our analysis by considering general oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes are being removed to form these partitions. This is a conceptual idea. Instead, Think of it as putting a box around the partition. Now the nodes inside the box are can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another, but thinking of them in this way, allows us to do more.
This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecure. Many of these approaxhes were conceptual or theoretical, saying only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper. Secondly, most of these paper focused on local search techniques that statrted at or near the desired degree doubling node. That presents a problem because oriented graphs can be varying in nature, This is the first exhaustive search technique on the SSNC problem.
We have laid the groundwork for our algorithm to tackle the oriented graphs. With a minimum degree node, we can partition this graph into independent subproblems. This is now ready for an exhaustive search algorithm.
@@ 4 Initial Lemmas
Minimum Out-Degree $<$ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
@@
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0. Hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. Since $v_{0}$ is a minimum out-degree node, we know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$. Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
@@
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, in order to construct our counter-example, we need out minimum degree to be at least 3. This will also place that same minimum value restriction on every other node in the graph. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
@@
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$. Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
@@
Thus, $v_{0}$’s out-degree will at least double in $G^{2}$. ∎
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$ shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
@@
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
@@
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) answers a question that we will come back to throughout this paper - this is an exhaustive search algorithm. In order for that to work, it has to allow some node’s out-degree to not double. In this case, what we see is that the initial node in the example $v_{0}$ does not have its degree double. Instead, we are forced to search for the next node. That comes at a neighbor of $v_{0}$, $v_{1}$ whose degree does double in this example.
This is an important distinction because we did not need the ”right” starting node. We chose a minimum degree node because of its ability to partition the nodes in the graph, not for its potential closeness to the double degree node.
For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph. This can be done because the nodes within a common induced subgraph are all relating to the same set of nodes.
These lemmas provide insight for our motivation, and a framework for theoretical foundation. What is coming in the next sections will build on these results and will refer back to them.
@@ 5 Theoretical Foundation
In order to understand this next section, we will introduce some new terms. We will give a formal name to the partitions we spoke of about previously - containers. We will also speak about the decreasing sequence property and talk about its importance. Then we will talk about interior, exterior and back arcs. Then we will close with the GLOVER data structure.
@@
Daamouch used induced subgraphs similar in his study of $m$-free graphs [7]. The primary differece here is the neighboring containers (see image) which will be important in the proofs.
With the concept of containers we can explore the structure of oriented graphs from two perspectives. We have the option of exploring it from the perspective of the minimum degree node, or from the perspective of containers.
  • Containers are seen based on their distance from the
  • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
minimum degree node.
@@
We have now introduced two concepts: the minimum degree node and the decreasing sequence property (DSP). A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point both for our algorithm and degree values. By selecting it as the starting point it, it serves as the anchor. Now, the minimum degree node is forcing the remaining nodes into the properties that will emerge in the counterexample construction. It is acting as a catalyst.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
@@
Unlike a simple total order, which could be any sequence with a single, consistent ordering (like lexicographic ordering), the container structure allows nodes at the same distance to be further organized. This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties.